OpenStudy (anonymous):

20 is 50% of what number? If necessary, round to the nearest tenth. Is it 10.0?

OpenStudy (ash2326):

100 % of a no. is that no. itself if the percent is less than 100 then the resultant no. will be less than our original no. Here 20 is 50% of some no. so that no. has to be greater than 20. Do you get this?

OpenStudy (anonymous):

oh okay :) so then it would be like 40?

OpenStudy (ash2326):

yeah, but let's solve to find that \[50 \% x=20\] \[\frac {50}{100} \times x=20\] Now solve this for x

OpenStudy (anonymous):

50/100=0.5...so then..20/0.5 equals...40!!! :O :D wait did i do that right?

OpenStudy (ash2326):

COrrect

OpenStudy (anonymous):

:D yay thank you! and thank you for the medals!